Let T be an equilateral triangle of height h. Let S be a square of side s. If T and S have the same area, then find h/s.
Answer by edjones(8007)

A=.5sh
A=s^2
s^2=.5sh
s=.5h
s/h=.5
h/s=1/.5
h/s=2
.
Ed